RC Update and Run Plan for Aug 2, 2020
- 1 RC Meeting Information
- 2 General updates
- 3 Current Run Plan
- 4 Notices in Effect for Shift Crew
- 5 Planned Accelerator Configuration Changes
- 6 Useful Links
RC Meeting Information
RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room
Sunday Aug 2 2020.
- Hall C in beam permit and ready; some tune beam in the Hall
- Got tune beam in the morning
- Ops is struggling with the beam size at the end of the C line
- They are trying a new optics solution
- As soon as they can send CW for beam centering we will proceed with that
- SoLID Cher test rocked tha House!!!! https://logbooks.jlab.org/entry/3819362 , https://logbooks.jlab.org/entry/3819431 , https://logbooks.jlab.org/entry/3819462 , https://logbooks.jlab.org/entry/3819492
Recent polarization measurements
Current Run Plan
First beam into the Hall
- Tune beam is allowed; make sure the target ladder is in "no target" position - MCC should notify the RC one hour prior to tune beam arrival in Hall C
- Before allowing any CW into Hall C we have to go through the beam centering procedure to establish optimal beam positions for the 3He cell and for the Reference cell to make sure there is no beam strike in the target region - expert driven (Hall C RC will call Brad when we are at that step)
- After the optimal beam positions are established MCC will take over the program to certify the machine tune; the target ladder must be in the "no target position" and the position lock must be engaged. The TO/SL must monitor that the beam positions stay at nominal values. We expect 30-40 muA to the dump as part of the "machine tune certification" exercise
- Once MCC will be ready to send Physics beam follow the plan outlined under Normal Production running :
Normal Production running
- Make sure operational restrictions are followed.
- Make sure the raster size is correct (100-500 mum)
- Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
- Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
- Change the IHWP every 4 hours when running on glass cells
- Run plan
If Physics beam is first time delivered during Swing or Owl:
- Check that the beam profile (most recent harp scans) and beam positions (orbit lock on) are O.K.; make sure the Hall C current ramp protection is enabled.
- Ion chamber calibrations on Reference cell filled with N2 @ 10 Atm (do not exceed a current of 30 muA)
- Ion chamber calibrations on Carbon Optics
- BCM calibration (you requested and MCC will follow the procedure)
- If there is still time before 8:00 am and the tatrget has good polarization then do production on 3He; if not Reference cell 10 atm Nitrogen (2h), Ref. cell 10 atm Hydrogen (2h), Ref. cell empty (2h)
- Make sure you reserve time to do an EPR calibration before the Moller measurement!!
If Physics beam is first time delivered during day shift:
- Dave G. will do a Moller measurement
Notes for tomorrow's 7:45 meeting
- After ~ 10 days of running Hall C will go to 1 pass for ~3 days
- None planned right now
When there is a down:
- If the down is 2+ hours call Simona: she will check with the RC to see if there is an opportunity for work in the Hall (SoLID Cherenkov parasitic test)
Notices in Effect for Shift Crew
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
- Spectrometer GUI (
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
Planned Accelerator Configuration Changes
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-220.127.116.11 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920